WILL YOU TEACH ME STEP BY STEP HOW TO SOLVE THE EQUATION (7X3+4X2+3X+2) + (2X2+X+5)
Answer by richwmiller(17219)

1) All caps is considered shouting on the internet.
2) There is no equation. Equations equate. Do you see an equal sign? I don't.
3) Use ^ for exponents such as x^2 x^3
(7x^3+4x^+3x+2) + (2x^2+x+5)
There is no need for parentheses here since it is just addition
7x^3+4x^2+3x+2 + 2x^2+x+5
so combine like terms
4x^2+2x^2=6x^2
3x+x=4x
2+5=7
7x^3+6x^2+4x+7